OpenStudy (megannicole51):

Find the integral of (x^6+4)^2 dx

OpenStudy (anonymous):

square it!

OpenStudy (anonymous):

\[\int(x^{12}+8x^6+16)dx\]

OpenStudy (megannicole51):

so....(x^6+4)(x^6+4)? and use foil to calculate?

OpenStudy (anonymous):

square, yeah

OpenStudy (megannicole51):

hahah thats funny!

OpenStudy (megannicole51):

so is that the answer? i feel like it can't be that simple!

OpenStudy (anonymous):

you made it to calculus right? time to drop "foil" from your vocab and no, that is not the answer, that is how you are going to find the answer right?

OpenStudy (anonymous):

integrate term by term

OpenStudy (dan815):

the foil bomb

OpenStudy (anonymous):

\[\int x^{12}dx+8\int x^6dx+\int 16dx\]

OpenStudy (megannicole51):

yeah im in calc 2 and at university....i like calling it foil.

OpenStudy (anonymous):

use the power rule backwards for each \[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

you good from there?

OpenStudy (dan815):

do they start integration in calc 2?

OpenStudy (megannicole51):

yeah they do

OpenStudy (megannicole51):

my professor doesnt speak very good english so I'm trying to learn the material and try to understand what hes saying lol

OpenStudy (megannicole51):

and what do u mean the power rule backwards?

OpenStudy (dan815):

watch youtube videos on it there are a lot

OpenStudy (dan815):

derivative backwards

OpenStudy (dan815):

just think what do i need to take the derivative of to get that

OpenStudy (dan815):

like 16... d/dx of 16x = 16

OpenStudy (dan815):

d/dx of 16x+c is also=16

OpenStudy (megannicole51):

ooooh i get it but how would u do like x^10?

OpenStudy (dan815):

for x^6 .... d/dx of 1/7*x^7 = x^6

OpenStudy (megannicole51):

so it would be x^9?

OpenStudy (anonymous):

\[\int x^{10}dx=\frac{x^{11}}{11}\] etc

OpenStudy (dan815):

d/dx 1/11 * x^11=x^10

OpenStudy (anonymous):

add one to the exponent, then divide by that number

OpenStudy (dan815):

u addin +1 to exponent and dividing by thaat exponent

OpenStudy (dan815):

the inverse of derivative power formula

OpenStudy (anonymous):

this is what i means by the power rule backwards \[\int x^ndx=\frac{x^{n+1}}{n+1}\] since the power rule is \[\frac{d}{dx}[x^n]=nx^{n-1}\]

OpenStudy (dan815):

for derivavtive u multiple my exp then subtract -1 from exp so backwards is +1 to exp and divide by exp

OpenStudy (dan815):

u multiply*

OpenStudy (anonymous):

power rule multiply by the exponent, drop the power by 1 power rule backwards add one to the exponent, divide by it

OpenStudy (dan815):

loser!! buddy

OpenStudy (dan815):

hows java coming along

OpenStudy (megannicole51):

so if i had the integral of (x^12 dx+8) that would just be x^11?

OpenStudy (dan815):

no

OpenStudy (megannicole51):

okay im confused im sorry

OpenStudy (dan815):

ok ill make it simple

OpenStudy (dan815):

derivative of x^2 =?

OpenStudy (megannicole51):

2x

OpenStudy (dan815):

integral of 2x =?

OpenStudy (megannicole51):

2?

OpenStudy (dan815):

no its x^2

OpenStudy (megannicole51):

okay

OpenStudy (dan815):

try this

OpenStudy (megannicole51):

okay:)

OpenStudy (dan815):

3x^2

OpenStudy (dan815):

integral of that?

OpenStudy (dan815):

its gonna be confusing for the first week u will mix up derivatives and integrals, we all did, but u will get used to it

OpenStudy (megannicole51):

okay i dont understand how you are taking the integral simple equations....i dont remember how to do them

OpenStudy (megannicole51):

thanks for the reassurance :)

OpenStudy (dan815):

okAY look carefully what we are doing for derivative

OpenStudy (megannicole51):

yes i can take the derivative

OpenStudy (dan815):

d/dx of x^3 =

OpenStudy (megannicole51):

3x^2

OpenStudy (megannicole51):

so the integral is x^3?

OpenStudy (anonymous):

\[\int xdx=\frac{x^2}{2}\\\int x^2dx=\frac{x^3}{3}\\\int x^3dx=\frac{x^4}{4}\\\int x^4dx=\frac{x^5}{5}\\\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (dan815):

yeah

OpenStudy (dan815):

break up the steps though, so for derivative u first multipled by the exponent and subtracted 1 from the exponent

OpenStudy (megannicole51):

so x^12 is x^13/13?

OpenStudy (dan815):

yes

OpenStudy (megannicole51):

ooooh okay

OpenStudy (megannicole51):

so what do u do if u just have say 5?

OpenStudy (dan815):

5x

OpenStudy (dan815):

theres a catch to all this though, there can be a constant there too

OpenStudy (dan815):

for example d/dx 5x+1 = 5 so is d/dx 5X +2 = 5 d/dx 5x+c = 5

OpenStudy (dan815):

so just keep doing whatever u are doing, but after u get ur integral answer just add a c to it

OpenStudy (megannicole51):

okay so for the actual problem it would be the integral of x^13/13+8x integral x^7/7 dx + integral 16x dx?

OpenStudy (megannicole51):

sorry i dont know how to make my equations look fancy lol

OpenStudy (dan815):

the integral and dx sighs are lost once u do the integration buy yes

OpenStudy (megannicole51):

so is that the answer to the problem? or is there more?

OpenStudy (dan815):

x^13/13+8x +x^7/7 +16x +c

OpenStudy (dan815):

that sthe answer

OpenStudy (dan815):

u can simplify it but ya

OpenStudy (megannicole51):

why is it set up like that? do u know?

OpenStudy (dan815):

like what

OpenStudy (megannicole51):

why do they take out the integrals?

OpenStudy (dan815):

its just an operator sign like d/dx

OpenStudy (dan815):

after u do d/dx of 5x

OpenStudy (dan815):

u will get 5 not d/dx of 5 right

OpenStudy (megannicole51):

my online thing says its wrong:(

OpenStudy (dan815):

samething with integral and dx d/dx mEANS DERIVATIVE WITH RESPECT TO X integral dx means integral with respect to x

OpenStudy (megannicole51):

so like change the 8x to 8? and 16x to 16?

OpenStudy (dan815):

what do u mean

OpenStudy (megannicole51):

the answer was wrong is there something we are missing?

OpenStudy (dan815):

did u add the c

OpenStudy (megannicole51):

yes

OpenStudy (dan815):

try without it then

OpenStudy (megannicole51):

nope didnt work

OpenStudy (dan815):

what are u entering

OpenStudy (megannicole51):

x^13/13+8x+x^7/7+16x+c

OpenStudy (dan815):

its 8x^7/7

OpenStudy (dan815):

why dont u add the 8x and 16x together too

OpenStudy (megannicole51):

then is it 16x^15/15?

OpenStudy (dan815):

wuuttt lool